# A stochastic network approach to clustering and visualising single-cell genomic count data

### v4

@@ Abstract
Important tasks in the study of genomic data include the identification of groups of similar cells (for example by clustering), and visualisation of data summaries (for example by dimensional reduction). In this paper, we propose a novel approach to studying single-cell genomic data, by modelling the observed genomic data count matrix $\mathbf{X}\in\mathbb{Z}_{\geq 0}^{p\times n}$ as a bipartite network with multi-edges. Utilising this first-principles network representation of the raw data, we propose clustering single cells in a suitably identified d-dimensional Laplacian Eigenspace (LE) via a Gaussian mixture model (GMM-LE), and employing UMAP to non-linearly project the LE to two dimensions for visualisation (UMAP-LE). This LE representation of the data estimates transformed latent positions (of genes and cells), under a latent position model of nodes in a bipartite stochastic network. We demonstrate how these estimated latent positions can enable fine-grained clustering and visualisation of single-cell genomic data, by application to data from three recent genomics studies in different biological contexts. In each data application, clusters of cells independently learned by our proposed methodology are found to correspond to cells expressing specific marker genes that were independently defined by domain experts. In this validation setting, our proposed clustering methodology outperforms the industry-standard for these data. Furthermore, we validate components of the LE decomposition of the data by contrasting healthy cells from normal and at-risk groups in a machine-learning model, thereby identifying an LE cancer biomarker that significantly predicts long-term patient survival outcome in two independent validation cohorts with data from 1904 and 1091 individuals.
@@ Introduction
Network models and methods have become an important subject in modern statistical science, especially in application-areas such as cell biology. A network can be represented by an ‘adjacency matrix’ [bose1963strongly], denoted by $\mathbf{A}$ (Figure LABEL:introFig). In the simplest case, a network consists of nodes of only one type (called a ‘unipartite network’), and these nodes may be connected (or not) by exactly one unweighted edge; in this case, adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$ (Figure LABEL:introFiga), where $n$ is the number of nodes in the network. When a network has nodes of two different types (represented by blue circles and red squares in Figure LABEL:introFigb), we call this a ‘bipartite’ network; an example of a corresponding asymmetric adjacency matrix is $\mathbf{A}\in\{0,1\}^{p\times n}$. Network models are well established in genomics, and a stochastic network model [erdHos1960evolution] refers to a statistical model of a network in which there is a probability $p_{ij}$ (defined according to some model) of observing an edge between nodes i and j. The edge-probability $p_{ij}$ is modelled according to a law which may depend on some data on or characteristics of the nodes i and j such as unobserved latent features $\phi_{i}^{(L)}$ and $\phi_{j}^{(R)}$ [hoff2002latent, young2007random], where we use the superscripts .$^{(L)}$ and .$^{(R)}$ to differentiate between vectors corresponding to the two different types of nodes in our bipartite setting.
@@
In genomics, network models are typically used to represent gene regulation in a gene regulatory network [bray1995protein], or related notions such as the gene co-expression network. Gene regulation is the phenomenon in which the protein encoded by one gene activates or represses the expression of another gene. When the expression of one gene regulates the expression of another gene in this way, this individual gene regulation is represented by an edge connecting the repressing and activating genes in the gene regulatory network. However, the focus of the work presented in this paper is not gene regulatory networks. Instead, we use a network model as a novel representation of single-cell genomic count data, and we draw on recent advances in stochastic network modelling to improve clustering and visualisation of the single cells in a genomic data set. Existing methods such as the industry-standard Seurat software package [satija2015spatial] also model single-cell genomic count data with an adjacency matrix, by using the k nearest neighbours method to summarise the data matrix as a binary symmetric adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$. In contrast, we propose a first-principles model of the data in which the raw count data themselves can be viewed directly as an asymmetric adjacency matrix $\mathbf{A}\in\mathbb{Z}_{\geq 0}^{p\times n}$ with entries $A_{ij}$ that can take 0 or any positive integer. In other words, in our setup we propose modelling the data directly as a bipartite multi-edge network (Figure LABEL:introFigb).
@@
Raw genomic sequencing data (RNA-seq, ATAC-seq, etc) may be represented as a matrix $\mathbf{X}\in\mathbb{Z}_{\geq 0}^{p\times n}$ of non-negative integer counts. These are counts of genomic sequencing reads, such that each entry $X_{ij}$ represents the number of observed copies of genomic fragment (e.g., mRNA transcript) $i\in\{1,...,p\}$ in sample (e.g., cell) $j\in\{1,...,n\}$. Modern single-cell sequencing data use UMIs (unique molecular identifiers), which avoid double-counting the same transcript (after amplification by PCR, polymerase chain reaction). Historically, it was considered necessary to model dropouts in single-cell sequencing data, which manifest in the data as zero-inflation. However, it has been shown more recently that it is no longer necessary to model dropouts or zero-inflation if these data use UMIs [townes2019feature]. For UMI data, under the model of [townes2019feature] the columns of $\mathbf{X}$ are modelled from first-principles as $n$ independent draws from a multinomial distribution. This contrasts with long-established methods for analysing RNA-seq data which use normalisation methods such as TPM (transcripts per million) and FPKM (fragments per kilobase per million), as part of standardised pipelines. Normalisation methods such as TPM and FPKM normalise the data by library size (i.e., $\sum_{i=1}^{p}X_{ij}$), and this may induce spurious correlations between transcripts [townes2019feature], as well as making it difficult to write down expressions for the true error distributions of the data after normalisation. In other words, after normalisation, we can’t say exactly what the true distribution of the normalised data is, and so rely on heuristics. There is currently no consensus in the genomics literature on best practice for normalising and modelling raw single-cell sequencing data [hafemeister2019normalization, lause2021analytic, choudhary2022comparison]. Instead, we take a unique view of these raw count data that is inspired by stochastic network theory. Our perspective allows data modelling based on a fundamental understanding of the distribution of the raw count data, and the summaries that follow.
@@
When the genomic data being analysed are the only available observations of some single cells in a study, an important task when analysing these data is to group cells together of similar types for identification or ‘phenotyping’, e.g., by clustering. One of the most successful and reliable methods for clustering single cells in genomic data is the so-called ‘Louvain clustering’ method [blondel2008fast], as implemented in the industry-standard Seurat software package [satija2015spatial]. Louvain clustering works by optimising an estimator called the ‘network modularity’, which is discussed in Section 2.3. Another important task when analysing single-cell data is visualisation of the data by projection into two dimensions, e.g., using non-linear methods such as UMAP (uniform manifold approximation and projection). The main aim of these projections is that data points which are close together in the original high dimensional space should remain close together in the two dimensional projection. The p features of data matrix $\mathbf{X}\in\mathbb{R}^{p\times n}$ usually contain much redundancy, and so to aid computational tractability and reduce noise [haber2017single], spectral methods may be used to reduce the dimension of the data to $\mathbf{V}^{\top}\in\mathbb{R}^{d\times n}$, where $d\ll p$, before using a non-linear method such as UMAP to generate the two dimensional projection $\mathbf{Y}\in\mathbb{R}^{2\times n}$ from $\mathbf{V}^{\top}$. It is natural to carry out both spectral clustering and non-linear projection using the same spectral decomposition $\mathbf{V}$ of the data-matrix $\mathbf{X}$.
@@
The structure of this papers is as follows. In Section 2, we present our novel first-principles representation of single-cell genomic count data, and our improved method of clustering single cells that we call ‘GMM-LE’ clustering, and related visualisations that we call ‘UMAP-LE’. Then in Section 3, we present the results of applying GMM-LE clustering and UMAP-LE visualisation to example data-sets relevant to human cortical development, to human embryonic development, and to a population at risk of breast cancer, and we identify and validate a machine-learning cancer biomarker from the LE representation of the data.
@@ Methods
The model specification that we propose here is based on the observation that a single-cell genomic data-matrix $\mathbf{X}\in\mathbb{Z}_{\geq 0}^{p\times n}$ has equivalent characteristics to the adjacency matrix $\mathbf{A}$ that represents a bipartite network with multi-edges (Figure LABEL:introFigb). By ‘multi-edges’ we mean that a pair of nodes may be connected by multiple edges, that is, $A_{ij}\in\mathbb{Z}_{\geq 0}$. By ‘bipartite’ we mean that the network can have nodes of two different types (row-nodes and column-nodes, represented by blue circles and red squares in Figure LABEL:introFigb), and in our setting we restrict this so that edges can only connect nodes of different types. Furthermore, both $\mathbf{A}$ and $\mathbf{X}$ tend to be very sparse (i.e., contain a large proportion of zeros: typically fewer than 10% of the matrix elements are non-zero in both cases). Hence, modelling $\mathbf{X}$ with $\mathbf{A}$ provides a mathematical justification for using clustering methodology designed for networks, as we propose here. We note that this adjacency matrix is $\mathbf{A}\in\mathbb{Z}_{\geq 0}^{p\times n}$, and our model is thus specified as $\mathbf{A}=\mathbf{X}$.
@@
Having specified our model of the data as $\mathbf{A}=\mathbf{X}\in\mathbb{Z}_{\geq 0}^{p\times n}$, following the latent position model of Hoff [hoff2002latent] (Eq.1), we assume that edges $A_{ij}$, $i\in\{1,...,p\}$, $j\in\{1,...,n\}$, in the bipartite network occur independently according to density $f$, which is a function of unobserved latent position vectors $\phi^{(L)}_{i}$ and $\phi^{(R)}_{j}$:
$A_{ij}\sim f\left(\phi^{(L)}_{i},\phi^{(R)}_{j}\right).$ (1)
We note that $A_{ij}$ is the number of edges between the $i^{\text{th}}$ row-node and the $j^{\text{th}}$ column-node in the network model, which also corresponds to $X_{ij}$, the number of counts for the $i^{\text{th}}$ gene in the $j^{\text{th}}$ cell in the data-set. We also note that for $A_{ij}\in\{0,1\}$, the usual form of Eq.1 would be $A_{ij}\sim\text{Bernoulli}(p_{ij})$, where $p_{ij}=\tilde{f}(\phi^{(L)}_{i},\phi^{(R)}_{j})$.
The random dot-product graph (RDPG) [young2007random] and generalised random dot-product graph (GRDPG) [rubin2022statistical] are linear latent position models. Under these models, the probability $p_{ij}$ of observing an edge between nodes $i$ and $j$, defined as $p_{ij}=\mathbb{E}(A_{ij}|\phi^{(L)}_{i},\phi^{(R)}_{j})$, specifically takes the form of an inner product and generalised inner product under the RDPG and GRDPG, respectively. Under these models, the latent position vectors $\phi^{(L)}_{i}$ and $\phi^{(R)}_{j}$ can be estimated consistently (up to an orthogonal transformation) [sussman2012consistent, lyzinski2014perfect, lyzinski2016community, rubin2022statistical] by spectral clustering of the adjacency matrix (known as adjacency spectral embedding, ASE), or on the Laplacian matrix, known as Laplacian spectral embedding (LSE). For genomic data, we choose to work with the Laplacian matrix as it is known to allow discovery of ‘affinity’ structure (in the embedded space), in contrast with the ASE which is preferred for discovery of ‘core-periphery’ structure [priebe2019two]. In our proposed modelling setup for genomic data, affinity structure corresponds to grouping genes with similar patterns of expression across cells (via transformations of $\phi^{(L)}_{i}$ as the rows of $\mathbf{U}$), or grouping cells with similar patterns of expression across genes (via transformations of $\phi^{(R)}_{j}$ as the rows of $\mathbf{V}$).
@@
We assume that cells with similar patterns of gene expression represent cells with similar phenotypes. Therefore, clustering cells (samples) in the space of the RSVs (denoted $\mathbf{V}$) provides a sensitive method for unsupervised learning of known and unknown cell-types from unlabelled single-cell genomic data. In other words, we cluster together samples $j$ that have similar (transformed) latent positions $\phi^{(R)}_{j}$. We optimise this unsupervised learning procedure by learning from recent advances in understanding of the latent geometry of the LE that follow from the theory of stochastic networks. Following the central limit theorem established for latent position vectors estimated via LSE [athreya2016limit, tang2018limit, rubin2022statistical], we carry out this clustering by fitting a Gaussian mixture model in the LE and refer to it as GMM-LE clustering. Before clustering, we also normalise the rows of $\mathbf{U}$ and $\mathbf{V}$ to unit length (so that the latent positions are projected onto a unit hypersphere), to help take account of nodes with very small degrees [ng2001spectral, rohe2011spectral, lei2015consistency]. This normalisation is carried out after calculating ‘leverage scores’ [qin2013regularized] for each $j\in\{1,..n\}$ as $||\mathbf{v}_{j}||_{2}$, for use downstream in the GMM-LE clustering (Section 2.4), where $\mathbf{v}_{j}$ is the $j^{\text{th}}$ row of $\mathbf{V}$. If necessary, we also inflate the degree distributions $\mathbf{d}_{L}$ and $\mathbf{d}_{R}$ by their medians to regularise the spectral clustering [chaudhuri2012spectral, qin2013regularized, amini2013pseudo], to improve computational identifiability in the presence of many low-degree nodes.
@@
We carry out spectral clustering by by fitting a Gaussian mixture model (GMM) in the LE, which we call GMM-LE clustering. We start with a latent position model (Eq.1) of our bipartite graph representation of the data $\mathbf{X}$, which we model as $\mathbf{A}=\mathbf{X}\in\mathbb{Z}_{\geq 0}^{p\times n}$, corresponding to an asymmetric adjacency matrix $\mathbf{A}$. Hence, our proposed methodology allows separate but related clustering (Section 2.4) for the two different types of nodes (corresponding to latent positions $\phi^{(L)}_{i}$, $i\in\{1,...,p\}$ and $\phi^{(R)}_{j}$, $j\in\{1,...,n\}$). This separate clustering of nodes of two types is referred to as co-clustering, or co-community detection [rohe2016co, bartlett2021co].
Our approach contrasts with other recent approaches such as [modell2022spectral], in which nodes of both types are included in one unified clustering procedure, by spectral embedding of a symmetric adjacency matrix that has all nodes of all types on both margins as $\left[\begin{array}[]{cc}\mathbf{0}&\mathbf{A}^{\top}\\ \mathbf{A}&\mathbf{0}\end{array}\right],$ where $\mathbf{A}\in\mathbb{Z}_{\geq 0}^{p\times n}$ is our asymmetric adjacency matrix $\mathbf{A}=\mathbf{X}$, and $\mathbf{0}$ is an appropriately-sized matrix of 0s. Such a matrix always has eigenvalues in pairs of equal magnitude and opposite sign. Furthermore this symmetric embedding leads to corresponding pairs of reflected eigenvectors (Lemma 5 of [modell2022spectral]), and negative eigenvalues are known to indicate disassortative features [khor2010concurrency], which in our context allow separation of the two different types of nodes (genes and cells). For example, consider the 2-node bipartite network represented (in our setting) by the adjacency matrix $A=[1]$, i.e., the corresponding data matrix $X=[1]$ represents 1 count of 1 gene in 1 cell. Then, the symmetric embedding of $A$ described above would be $\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right],$ which has eigenvalues 1 and -1, with corresponding eigenvectors $[1/\sqrt{2},1/\sqrt{2}]^{\top}$ and $[-1/\sqrt{2},1/\sqrt{2}]^{\top}$. The second of these eigenvectors is disassortative, because dividing the nodes based on a separating point anywhere in $\left(-1/\sqrt{2},1/\sqrt{2}\right)$ will separate the nodes into two groups according to their types ‘cell’ and ‘gene’ as a ‘cut’ based on this second eigenvector [shi2000normalized, von2007tutorial]. Whereas there is no separating point for which the first eigenvector can be used to separate the nodes into groups.

### v5

@@ Abstract
Important tasks in the study of genomic data include the identification of groups of similar cells (for example by clustering), and visualisation of data summaries (for example by dimensional reduction). In this paper, we develop a novel approach to these tasks in the context of single-cell genomic data. To do so, we propose to model the observed genomic data count matrix $\mathbf{X}\in\mathbb{Z}_{\geq 0}^{p\times n}$, by representing these measurements as a bipartite network with multi-edges. Utilising this first-principles network model of the raw data, we cluster single cells in a suitably identified d-dimensional Laplacian Eigenspace (LE) via a Gaussian mixture model (GMM-LE), and employ UMAP to non-linearly project the LE to two dimensions for visualisation (UMAP-LE). This LE representation of the data-points estimates transformed latent positions (of genes and cells), under a latent position statistical model of nodes in a bipartite stochastic network. We demonstrate how transformations of these estimated latent positions can enable fine-grained clustering and visualisation of single-cell genomic data, by application to data from three recent genomics studies in different biological contexts. In each data application, clusters of cells independently learned by our proposed methodology are found to correspond to cells expressing specific marker genes that were independently defined by domain experts. In this validation setting, our proposed clustering methodology outperforms the industry-standard for these data. Furthermore, we validate components of the LE decomposition of the data by contrasting healthy cells from normal and at-risk groups in a machine-learning model, thereby identifying an LE cancer biomarker that significantly predicts long-term patient survival outcome in two independent validation cohorts with data from 1904 and 1091 individuals.
@@ Introduction
Network models and methods have become an important subject in modern statistical science, especially in application-areas such as cell biology. A network can be represented by an ‘adjacency matrix’, denoted by $\mathbf{A}$ (Figure LABEL:introFig). In the simplest case, a network consists of nodes of only one type (called a ‘unipartite network’), and these nodes may be connected (or not) by exactly one unweighted edge; in this case, adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$ (Figure LABEL:introFiga), where $n$ is the number of nodes in the network. When a network has nodes of two different types (represented by blue circles and red squares in Figure LABEL:introFigb), we call this a ‘bipartite’ network; an example of a corresponding asymmetric adjacency matrix is $\mathbf{A}\in\{0,1\}^{p\times n}$. Network models are well established in genomics, and a stochastic network model [erdHos1960evolution] refers to a statistical model of a network in which there is a probability $p_{ij}$ (defined according to some model) of observing an edge between nodes i and j. The edge-probability $p_{ij}$ is modelled according to a law which may depend on some data on or characteristics of the nodes i and j such as unobserved latent features $\phi_{i}^{(L)}$ and $\phi_{j}^{(R)}$ [hoff2002latent, young2007random] (i.e., a latent position model), where we use the superscripts $\cdot^{(L)}$ and $\cdot^{(R)}$ to differentiate between vectors corresponding to the two different types of nodes in our bipartite setting. In this manuscript we focus on methodology that follows a latent position model, although we note that alternative stochastic network models exist, such as the exponential random graph model (ERGM) [frank1986markov].
@@
In genomics, graphical models are often used to represent gene regulation in a gene regulatory network [bray1995protein], or related notions such as the gene co-expression network [dobra2004sparse, kolar2010estimating]. Gene regulation is the phenomenon in which the protein encoded by one gene activates or represses the expression of another gene. When the expression of one gene regulates the expression of another gene in this way, this individual gene regulation is represented by an edge connecting the repressing and activating genes in the gene regulatory network. In contrast, the focus of the work presented in this paper is not graphical modelling of gene regulatory networks. It is important to make clear that in contrast to that literature, we use a mathematical object viz. a graph and its corresponding adjacency matrix directly as a mathematical-biological model [murray2007mathematical] of single-cell genomic count data, in which numbers of mRNA molecules in specific cells are modelled by network edges. Then, we draw on recent advances in stochastic network theory in the mathematical-statistics literature to improve clustering and visualisation of the single cells in genomic data sets, based on our network representation of the data. Existing methods such as the industry-standard Seurat software package [satija2015spatial] also model single-cell genomic count data with an adjacency matrix, by using the k-nearest-neighbours method to summarise the data matrix as a binary symmetric adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$. In contrast, we propose a first-principles representation of the data in which the raw data-counts of mRNA molecules in specific cells can be modelled directly with an asymmetric adjacency matrix $\mathbf{A}\in\mathbb{Z}_{\geq 0}^{p\times n}$ with entries $A_{ij}$ that can take 0 or any positive integer. In other words, in our setup we propose modelling the data directly as a bipartite multi-edge network (Figure LABEL:introFigb).
@@
Raw genomic sequencing data (RNA-seq, ATAC-seq, etc) may be represented as a matrix $\mathbf{X}\in\mathbb{Z}_{\geq 0}^{p\times n}$ of non-negative integer counts. These are counts of genomic sequencing reads, such that each entry $X_{ij}$ represents the number of observed copies of genomic fragment (e.g., mRNA transcript) $i\in\{1,...,p\}$ in sample (e.g., cell) $j\in\{1,...,n\}$. Modern single-cell sequencing data use UMIs (unique molecular identifiers), which avoid double-counting the same transcript (after amplification by PCR, polymerase chain reaction). Historically, it was considered necessary to model dropouts in single-cell sequencing data, which manifest in the data as zero-inflation. However, it has been shown more recently that it is no longer necessary to model dropouts or zero-inflation if these data use UMIs [townes2019feature]. For UMI data, under the model of [townes2019feature] the columns of $\mathbf{X}$ are modelled from first-principles as $n$ independent draws from a multinomial distribution. This contrasts with long-established methods for analysing RNA-seq data which use normalisation methods such as TPM (transcripts per million) and FPKM (fragments per kilobase per million), as part of standardised pipelines. Normalisation methods such as TPM and FPKM normalise the data by library size (i.e., $\sum_{i=1}^{p}X_{ij}$), and this may induce spurious correlations between transcripts [townes2019feature], as well as making it difficult to write down expressions for the true error distributions of the data after normalisation. In other words, after normalisation, we can’t say exactly what the true distribution of the normalised data is, and so rely on heuristics. There is currently no consensus in the genomics literature on best practice for normalising and modelling raw single-cell sequencing data [hafemeister2019normalization, lause2021analytic, choudhary2022comparison]. Our proposed approach enables data modelling based on a first-principles representation of these data. This allows important insights into the distributions of representations of these data-points, such as asymptotic multivariate Gaussianity in spectrally-embedded space.
@@
When the genomic data being analysed are the only available observations of some single cells in a study, an important task when analysing these data is to group cells together of similar types for identification or ‘phenotyping’, e.g., by clustering. One of the most successful and reliable methods for clustering single cells in genomic data is the so-called ‘Louvain clustering’ method [blondel2008fast], as implemented in the industry-standard Seurat software package [satija2015spatial]. Seurat-Louvain clustering works by finding a k-nearest-neighbour graph and then optimising an estimator called the ‘network modularity’, which is discussed in Section 2.3. Another important task when analysing single-cell data is visualisation of the data by projection into two dimensions, e.g., using non-linear methods such as UMAP (uniform manifold approximation and projection). Whilst care must be taken with interpretation of UMAP (and alternative non-linear) embeddings [bohm2022attraction], the main aim of these projections is that data points which are close together in the original high dimensional space should remain close together in the two dimensional projection. It follows from the foundational model of modern biology (the ‘central dogma of biology’ [crick1958protein]), that cells with similar gene expression profiles are considered to have similar phenotypes. So while directions may become distorted under projections such as UMAP, in general we can expect cells with similar phenotypes to remain located close together in the two dimensional projection. This means that UMAP projections of genomic data are very useful for visual validation and identification of clusters corresponding to specific cellular phenotypes. The p features of data matrix $\mathbf{X}\in\mathbb{R}^{p\times n}$ usually contain much redundancy, and so to aid computational tractability and reduce noise [haber2017single], spectral methods may be used to reduce the dimension of the data to $\mathbf{V}^{\top}\in\mathbb{R}^{d\times n}$, where $d\ll p$, before using a non-linear method such as UMAP to generate the two dimensional projection $\mathbf{Y}\in\mathbb{R}^{2\times n}$ from $\mathbf{V}^{\top}$ [modell2021graph]. It is natural to carry out both spectral clustering and non-linear projection using the same spectral decomposition $\mathbf{V}$ of the data-matrix $\mathbf{X}$.
@@
The structure of this papers is as follows. In Section 2, we present our novel first-principles model of single-cell genomic count data, and based on this our improved method of clustering single cells that we call ‘GMM-LE’ clustering, and related visualisations that we call ‘UMAP-LE’. Then in Section 3, we present the results of applying GMM-LE clustering and UMAP-LE visualisation to example data-sets relevant to human cortical development, to human embryonic development, and to a population at risk of breast cancer. Finally, we identify and validate a machine-learning cancer biomarker obtained in a novel way from the LE representation of the data.
@@ Methods
The mathematical-biological model specification that we propose here follows from the observation that a single-cell genomic data-matrix $\mathbf{X}\in\mathbb{Z}_{\geq 0}^{p\times n}$ has equivalent characteristics to the adjacency matrix $\mathbf{A}$ that represents a bipartite network with multi-edges (Figure LABEL:introFigb). By ‘multi-edges’ we mean that a pair of nodes may be connected by multiple edges, that is, $A_{ij}\in\mathbb{Z}_{\geq 0}$. By ‘bipartite’ we mean that the network can have nodes of two different types (row-nodes and column-nodes, represented by blue circles and red squares in Figure LABEL:introFigb), and in our setting we restrict this so that edges can only connect nodes of different types. Furthermore, both $\mathbf{A}$ and $\mathbf{X}$ tend to be very sparse (i.e., contain a large proportion of zeros: typically fewer than 10% of the matrix elements are non-zero in both cases). Hence, we propose a mathematical framework where the numbers of mRNA molecules in specific cells are represented by a mathematical object viz. a graph and its corresponding adjacency matrix, i.e. $\mathbf{A}\in\mathbb{Z}_{\geq 0}^{p\times n}$. Representing $\mathbf{X}$ by $\mathbf{A}$ hence provides a mathematical justification for using clustering methodology informed by stochastic network modelling, as follows.
@@
Having specified our representation of the genomic data-set as $\mathbf{A}=\mathbf{X}\in\mathbb{Z}_{\geq 0}^{p\times n}$, we assume that the $A_{ij}$ in the corresponding bipartite network occur independently according to some probability mass function (PMF) $p_{A_{ij}}\in[0,1]$ as
$\text{Pr}\left(A_{ij}\right)=p_{A_{ij}}\left(A_{ij};\phi^{(L)}_{i},\phi^{(R)}_{j}\right);\enspace\sum_{A_{ij}=0}^{\infty}p_{A_{ij}}\left(A_{ij};\phi^{(L)}_{i},\phi^{(R)}_{j}\right)=1,\enspace i\in\{1,...,p\},\enspace j\in\{1,...,n\},\vspace*{1ex}$ (1)
such that $E(A_{ij}|\phi^{(L)}_{i},\phi^{(R)}_{j})$ admits a low-rank representation. We note that in our setting, $A_{ij}$ is the number of edges between the $i^{\text{th}}$ row-node and the $j^{\text{th}}$ column-node in the network model, which corresponds to $X_{ij}$, the number of counts for the $i^{\text{th}}$ gene in the $j^{\text{th}}$ cell in the genomic data-set. Genomic count data are typically over-dispersed compared to a Poisson model [townes2019feature, hafemeister2019normalization, lause2021analytic, choudhary2022comparison], and hence we work under the more general setting of Eq.1.
A canonical example of the latent position model of Hoff [hoff2002latent] with $A_{ij}\in\{0,1\}$ is $A_{ij}\sim\text{Bernoulli}(p_{ij})$ (which we note may be approximated by $A_{ij}\sim\text{Poisson}(p_{ij})$ [perry2012null]). In that setting, our model would reduce to $E(A_{ij}|\phi_{i},\phi_{j})=p_{ij}=\text{Pr}\left(A_{ij}=1|\phi_{i},\phi_{j}\right)$, which admits a low rank representation as often described in the literature under the random dot-product graph (RDPG) [young2007random] and generalised random dot-product graph (GRDPG) [rubin2022statistical] models, where $p_{ij}$ takes the form of an inner product and generalised inner product of latent position vectors, respectively. Under the (G)RDPG models, the latent positions can be estimated by the eigenvalues of the adjacency matrix scaled by the square-roots of the absolute values of their corresponding eigenvalues. More generally, under these models the latent position vectors $\phi^{(L)}_{i}$ and $\phi^{(R)}_{j}$ can be estimated consistently (up to an orthogonal transformation) [sussman2012consistent, lyzinski2014perfect, lyzinski2016community, rubin2022statistical] by spectral clustering of the adjacency matrix (known as adjacency spectral embedding, ASE) or of the Laplacian matrix, known as Laplacian spectral embedding (LSE). In this context, a Gaussian mixture model is appropriate to carry out the spectral clustering of the nodes, because their latent positions in the ASE, or the transformations of these latent positions in the LSE, are known to follow multivariate Gaussian distributions [rubin2022statistical]. In the LSE, the latent positions of the nodes are transformed as a result of the normalisation by the degree distributions of the nodes (or functions thereof) that takes place in the calculation of the graph Laplacian. Furthermore, while scaling by the square-root of the eigenvalue magnitudes is required to obtain the latent positions under the RDPG and GRDPG, such scaling will not change the asymptotic Gaussian assumption, because the effect of this scaling is to stretch the axes of the multivariate Gaussian distribution. It is also known that this asymptotic Gaussian assumption extends to weighted graphs [gallagher2023spectral], such as the multi-edge graphs that are the focus of this article. We therefore assume that the transformed latent positions of the nodes are asymptotically multivariate-Gaussian distributed in the LSE of the multi-edge graphs that we discuss here, as a subset of weighted graphs (i.e., with weights taking only integer values). Hence for the unsupervised learning and visualisation applications described in this article, we do not need to specify the precise form of the PMF of Eq.1, such as the overdispersion of the edge-counts $A_{ij}$.
@@
We assume that cells with similar patterns of gene expression represent cells with similar phenotypes. Therefore, clustering cells (samples) in the space of the RSVs (denoted $\mathbf{V}$) provides a sensitive method for unsupervised learning of known and unknown cell-types from unlabelled single-cell genomic data. In other words, we cluster together samples $j$ that have similar (transformed) latent positions $\phi^{(R)}_{j}$. We optimise this unsupervised learning procedure by following recent advances in understanding of the latent geometry of the LE that follow from the theory of stochastic networks. Following the central limit theorem established for (transformations of) latent position vectors estimated via LSE [athreya2016limit, tang2018limit, rubin2022statistical], we carry out this clustering by fitting a Gaussian mixture model in the LE and refer to it as GMM-LE clustering. Before clustering, we also normalise the rows of $\mathbf{U}$ and $\mathbf{V}$ to unit length (so that the transformed latent positions are projected onto a unit hypersphere), to help take account of nodes with very small degrees [ng2001spectral, rohe2011spectral, lei2015consistency]. This normalisation is carried out after calculating ‘leverage scores’ [qin2013regularized] for each $j\in\{1,..n\}$ as $||\mathbf{v}_{j}||_{2}$, for use downstream in the GMM-LE clustering (Section 2.4), where $\mathbf{v}_{j}$ is the $j^{\text{th}}$ row of $\mathbf{V}$. If necessary, we also inflate the degree distributions $\mathbf{d}_{L}$ and $\mathbf{d}_{R}$ by their medians to regularise the spectral clustering [chaudhuri2012spectral, qin2013regularized, amini2013pseudo], to improve computational identifiability in the presence of many low-degree nodes.
@@
We carry out spectral clustering of the raw genomic count matrix by fitting a Gaussian mixture model (GMM) in the LE, which we call GMM-LE clustering. We start with a latent position model (Eq.1) of our bipartite graph representation of the count-data matrix $\mathbf{X}$, i.e. $\mathbf{A}=\mathbf{X}\in\mathbb{Z}_{\geq 0}^{p\times n}$, noting the asymmetric adjacency matrix $\mathbf{A}$. Hence, our proposed methodology allows separate but closely connected clustering and analysis (Section 2.4) of the two different types of nodes (corresponding to latent positions $\phi^{(L)}_{i}$, $i\in\{1,...,p\}$ and $\phi^{(R)}_{j}$, $j\in\{1,...,n\}$). This separate clustering of nodes of two types is referred to as co-clustering, or co-community detection [rohe2016co, bartlett2021co]. Our approach is related to other recent approaches such as [modell2022spectral], in which a bipartite graph refers to one in which nodes may be partitioned into two different intrinsic groups. Those groups correspond to the two different types of nodes in our setup, which can be identified equivalently by spectral embedding of a symmetric adjacency matrix that has all nodes of all types on both margins, i.e. $\left[\begin{array}[]{cc}\mathbf{0}&\mathbf{A}^{\top}\\ \mathbf{A}&\mathbf{0}\end{array}\right],$ where $\mathbf{A}\in\mathbb{Z}_{\geq 0}^{p\times n}$ is our asymmetric adjacency matrix $\mathbf{A}=\mathbf{X}$, and $\mathbf{0}$ is an appropriately-sized matrix of 0s. Such a matrix always has eigenvalues in pairs of equal magnitude and opposite sign, and negative eigenvalues are known to indicate disassortative features [khor2010concurrency], which in our context allow separation of the two different types of nodes (genes and cells). For example, consider the 2-node bipartite network represented (in our setting) by the adjacency matrix $A=[1]$, i.e., the corresponding data matrix $X=[1]$ represents 1 count of 1 gene in 1 cell. Then, the symmetric embedding of $A$ described above would be $\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right],$ which has eigenvalues 1 and -1, with corresponding eigenvectors $[1/\sqrt{2},1/\sqrt{2}]^{\top}$ and $[-1/\sqrt{2},1/\sqrt{2}]^{\top}$. The second of these eigenvectors is disassortative, because dividing the nodes based on a separating point anywhere in $\left(-1/\sqrt{2},1/\sqrt{2}\right)$ will separate the nodes into two groups according to their types ‘cell’ and ‘gene’ as a ‘cut’ based on this second eigenvector [shi2000normalized, von2007tutorial]. Whereas there is no separating point for which the first eigenvector can be used to separate the nodes into groups.
@@
For example, consider the SVD decompositions $A_{ij}=\sum_{l=1}^{d}U_{i,l}S_{l,l}V_{l,j}$ and $A_{ij^{\prime}}=\sum_{l=1}^{d}U_{i,l}S_{l,l}V_{l,j^{\prime}}$ (for nodes $j$ and $j^{\prime}$). We wish to cluster together cells (represented by transformations of the latent positions modelled by the rows of $\mathbf{V}$) if those cells have large counts of transcripts from similar sets of genes (i.e., similar gene expression profiles), indicating similar phenotypes. Then if $A_{ij}$ represents the number of genomic counts $X_{ij}$, consider the contribution to counts $X_{ij}$ and $X_{ij^{\prime}}$ from SVD component $l$ in cases 1 and 2 below. Here we define gene $i$ as a marker gene for a certain cell type, $X_{ij}$ and $X_{ij^{\prime}}$ as the counts for marker gene $i$ in cells $j$ and $j^{\prime}$, and assume $U_{i,l}>0$: